OpenStudy (anonymous):

type an ordered pair....x+4y=9, x=3-4y

OpenStudy (anonymous):

what do u mean by type... do you want to solve for x and y?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1. Get them into y=mx=b form 1st equation= y=-1/4 x+2.25 2nd equation= y= 1/4 x -.75 p.s. it didn't have 2 be y=mx+b, the equations just had to be in the same style. 2. This might be confusing if you've never done it before, but you can just subtract one equation from the other. The point is to get rid of one variable so you can just concentrate on the other. since the y's are the same, you don't have to worry about doing any fancy multiplication work. Just subtract the second equation from the first. And remember to subtract everything- not just the y. 0= -1/2 x +3 It's easy to solve from there.

OpenStudy (anonymous):

I forgot part two- after you find out what x is, plug it back into one of the original equations and then solve for y. To check you're work, plug your values into the second equation and see if it works out.